An annual dividend of $22.58 for 2019 has just been announced by the CEO of That’s The Way You Live Your Life, D.S. Sicko. The firm is paying out 43% of the its 2019 earnings. D.S. Sicko announces that firm has earned a return of 11% on the retained earnings it reinvested in new projects since last year’s 2018 dividend, where it paid out 63% of that year’s earnings. Which of the following is closest to the dividend that was paid out in 2018?

a. $33.40 b. $23.58 c. $31.13 d. $31.79 e. $32.68
Answer #1

Answer: d. 31.79

Total earnings this year = $22.58÷0.43 = 52.51

The extra % it earned from over last year = Last year retained earnings×return on Investment.

Last year retained earnings = (1-0.63)×11 = 4%

Earning last year = 52.51/1.04
= 50.49

dividend that was paid out in 2018 = 50.49×63% = $31.80